Corrections to “Random Lazy Random Walks on Arbitrary Finite Groups” by Martin Hildebrand J. Theoret. Probab. 14 (2001), 1019-1034.

On the second line from the bottom, it is claimed that x is a function of g_{1}, …, g_{j}. In fact, x is a function of g_{1}, …, g_{ℓ} (as described in Proposition 7). The author regrets not detecting the incorrect subscript before publication.

The author also wishes to acknowledge that the concerns raised on p. 1026 about a proof of Pak were previously raised by the author in a Mathematical Reviews review of Pak’s paper. The author hopes that this omission of acknowledgment in the paper does not mislead readers.